9. A good’s demand is given by P =400-2Q At P=80 the point elasticity is
Answer #1

P= 400 - 2Q

2Q= 400-P

Q= 200 - 0.5P

dQ/dP = -0.5

At P= 80, Q= 200-0.5*80= 160

Elasticity = dQ/dP * P/Q = -0.5*80/160= -0.25